OpenStudy (anonymous):

help Use the three steps to solve the problem. One number is five more than another and their sum is three less than three times the smaller. Find the numbers.

OpenStudy (texaschic101):

x = y + 5 (one number is 5 more then another number) x + y = 3y - 3 (their sum (x + y) is (=) 3 less then 3 times the smaller (3y - 3...y being the smaller number) sub in y + 5 in for x (y + 5) + y = 3y - 3 2y + 5 = 3y - 3 2y - 3y = -3 - 5 -y = -8 y = 8 now sub 8 in for y x = y + 5 x = 8 + 5 x = 13 check... x + y = 3y - 3 13 + 8 = 3(8) - 3 21 = 24 - 3 21 = 21 (correct) larger number (x) = 13 smaller number (y) = 8

OpenStudy (anonymous):

thank u...;)

OpenStudy (texaschic101):

no problem :)